ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

5. Differential Geometry and Global Analysis

1991 MS Classification: Li, Congming, Univ. of Colorado. Prescribing Scalar Curvature on S^{n} The following two results on prescribing scalar curvature will be presented. 1: A necessary and sufficient condition for the class of rotationally symmetric functions. 2: A prior estimates for the solutions.